# Slot Attention with Re-Initialization and Self-Distillation

### v1

@@ Abstract.
Unlike popular solutions based on dense feature maps, Object-Centric Learning (OCL) represents visual scenes as sub-symbolic object-level feature vectors, termed slots, which are highly versatile for tasks involving visual modalities. OCL typically aggregates object superpixels into slots by iteratively applying competitive cross attention, known as Slot Attention, with the slots as the query. However, once initialized, these slots are reused naively, causing redundant slots to compete with informative ones for representing objects. This often results in objects being erroneously segmented into parts. Additionally, mainstream methods derive supervision signals solely from decoding slots into the input’s reconstruction, overlooking potential supervision based on internal information. To address these issues, we propose Slot Attention with re-Initialization and self-Distillation (DIAS): i) We reduce redundancy in the aggregated slots and re-initialize extra aggregation to update the remaining slots; ii) We drive the bad attention map at the first aggregation iteration to approximate the good at the last iteration to enable self-distillation. Experiments demonstrate that DIAS achieves state-of-the-art on OCL tasks like object discovery and recognition, while also improving advanced visual prediction and reasoning. Our code is available on https://github.com/Genera1Z/DIAS.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755339isbn: 979-8-4007-2035-2/2025/10submissionid: mfp3301copyright: noneccs: Computing methodologies Computer vision representations
@@
Figure 1. Two key observations inspire our method. (o1) We re-initialize an extra aggregation to update the remaining slots after slots redundancy reduction, instead of decoding the remaining slots directly. (o2) We self-distill the attention map at the first iteration to approximate the almost-always-good attention at the last aggregation iteration, rather than from aggregation attention to decoding attention.
@@ 1. Introduction
Humans’ vision system typically decomposes scenes into objects and allocates attention at the object level (Chen, 2012; Kahneman et al., 1992). This efficiently captures scene information, models object attributes and their interrelationships, and ultimately supports complex functionalities of prediction, reasoning, planning and decision-making (Bar, 2004). Similarly, in Artificial Intelligence (AI), Object-Centric Learning (OCL) seeks to aggregate dense (super-)pixels of images or videos into sparse feature vectors, termed slots, representing corresponding objects in a sub-symbolic manner (Locatello et al., 2020). In multimodal AI systems, object-centric representations can encode visual information more effectively and integrate with other modalities more flexibly than popular approaches based on dense feature maps (Zhang et al., 2024; Wang et al., 2024).
@@
Existing OCL typically follows the encoding-aggregation-decoding paradigm (Zhao et al., 2025b). Firstly, input image or video pixels are encoded into feature map; Secondly, object super-pixels in the feature map are iteratively aggregated into corresponding feature vectors, i.e., slots, using competitive cross attention, i.e., Slot Attention (Locatello et al., 2020) or its variants (Biza et al., 2023; Jia et al., 2023); Lastly, these slots are decoded to reconstruct the input in some format, providing the self-supervision signal.
@@
However, (i1) once initialized, these slots are iterated naively in the aggregation, and when there are fewer objects than the predefined number of slots, redundant slots would compete with informative ones for object representation (Aydemir et al., 2023; Fan et al., 2024), causing objects to be segmented into parts. Additionally, (i2) mainstream methods (Singh et al., 2022a; Seitzer et al., 2023; Wu et al., 2023b) derive supervision signals solely from reconstructing the input using the aggregated slots, leaving internal information unutilized, e.g., the aggregation attention maps at latter iterations have better object segmentation, which can be utilized to provide extra supervision signal.
@@
Issues (i1) and (i2) remain unresolved despite just a few ingenious attempts. (r1) SOLV (Aydemir et al., 2023) and AdaSlot (Fan et al., 2024) remove possible redundancy in the aggregated slots, but without further updating the remaining. As shown in Figure 1 o1, the remaining slots, despite being informative, are already distracted by redundant slots and are still of low quality in object representation. (r2) SPOT (Kakogeorgiou et al., 2024) drives the aggregation attention map to approximate the whole decoding attention map, but without discarding the latter’s bad compositions. As shown in Figure 1 o2, the decoding attention map, despite being generally better, sometimes has lower accuracy in object segmentation than the aggregation attention map. This is why SPOT has to pretrain an extra copy of the model as the teacher and has to run both the teacher and student for their offline distillation, at least doubling the training cost.
@@
Our solution is simple yet effective. (s1) We use the remaining slots after slots redundancy reduction to re-initialize an extra aggregation. This improves slots’ object representation quality, instead of directly decoding the remaining slots. (s2) We use the attention map at the last aggregation iteration to guide that at the first iteration, where the guidance is mostly always better. This enables self-distillation, requiring no extra model copy, no pretraining, and minimal additional cost. Besides, (s3) we implement a generalized auto-regressive (AR) decoder for OCL. Since existing AR decoders’ fix-order flattening disrupts spatial correlations among superpixels (Singh et al., 2022a, b; Kakogeorgiou et al., 2024), we use random-order flattening to enforce the decoder to model spatial correlations.
@@
Figure 2. Our DIAS introduces three novel designs: (i) Reinitialized aggregation, which improves slots’ object representation quality by re-initializing an extra aggregation to update the remaining slots after slots redundancy reduction; (ii) Self-distilled aggregation, which obtains better internal supervision by approximating the almost-always-better attention map at the last aggregation iteration from that at the first aggregation iteration; and (iii) Random auto-regressive decoding, which enforces the decoder’s modeling of spatial correlations by randomly flattening a 2-dimensional feature map into a 1-dimensional sequence. On the top left is the common OCL architecture, which is adapted from (Zhao et al., 2025b). $\bm{\phi}_{\mathrm{a}}$ is OCL aggregator, $\bm{\phi}_{\mathrm{d}}$ is OCL decoder, and $\bm{\phi}_{\mathrm{r}}$ is slots redundancy reduction. m is the mask token, bos is the Begin-of-Sentence token, and pos emb stands for the position embedding tensor. All notations are defined in the marked sections.
@@ 2. Relate Work
OCL aggregation. The core module of OCL is Slot Attention (Locatello et al., 2020), an iterative cross attention that takes some feature vectors as the competitive query, with the input’s feature map as the key and value. BO-QSA (Jia et al., 2023) improves the gradient flow on the query vectors for better optimization. ISA (Biza et al., 2023) realizes scaling, rotation and translation invariance in slots with inductive biases. While others learn Gaussian distributions to sample the query vectors, SAVi (Kipf et al., 2022) and SAVi++ (Elsayed et al., 2022) project conditions like object bounding boxes into query vectors. These aggregators are all faced with the issue that redundant slots compete with informative slots to represent single objects as parts. SOLV (Aydemir et al., 2023) merges similar slots after aggregation using agglomerative clustering by some threshold. AdaSlot (Fan et al., 2024) removes redundant slots after aggregation using a learned policy with some regularization. These two recent methods only remove or merge slots but do not update the remaining ones. Our DIAS improves the aggregation by both reducing redundant slots and updating the remaining slots with extra re-initialized aggregation.
@@
OCL decoding. There are mainly three types of decoding for OCL. Mixture-based decoders include CNN (Kipf et al., 2022; Elsayed et al., 2022), MLP (Seitzer et al., 2023) and SlotMixer (Zadaianchuk et al., 2024). The former two broadcast each of the aggregated slots into some spatial dimension, decode them respectively and mix the final components as the reconstruction. This means that the computation is proportional to the number of slots (Seitzer et al., 2023), which is very expensive. The latter one mixes the slots with a Transformer decoder without self-attention on the query. Such a decoder has to work in lower slot channel dimensions (Zadaianchuk et al., 2024), which limits object representation quality. Auto-regression-based decoders include causal Transformer decoder (Singh et al., 2022a, b) and causal Transformer decoder with nine permutations (Kakogeorgiou et al., 2024). They basically drive the slots to reconstruct the input’s feature map with the causally masked input feature as the query, in a next-token prediction manner. However, flattening a 2-dimensional feature map into a 1-dimensional sequence in a fixed order loses the spatial relationship (Kakogeorgiou et al., 2024). Diffusion-based decoders include different conditional Diffusion models (Wu et al., 2023b; Jiang et al., 2023). They basically recover the noise from the noise diffused input with slots as conditions, to drive slots to be informative. But, Diffusion models are very large and sensitive to train (Rombach et al., 2022). Our DIAS realizes general auto-regression via arbitrary ordering. In terms of the reconstruction target, there are also works that introduce quantization (Singh et al., 2022a, b; Zhao et al., 2024, 2025a, 2025b), but we focus on the non-quantized case.
@@
Knowledge distillation. The classical offline knowledge distillation (KD) (Gou et al., 2021) requires a pretrained model as the teacher and a to-be-trained model as the student. The teacher is frozen and runs at the same time as the student to provide better representations as targets for the corresponding student representations. There is also online distillation (Gou et al., 2021) where no pretraining of a teacher model is needed. A teacher model typically has the same architecture as the student and its weights are usually updated by exponentially moving averaging (Wikipedia, [n. d.]) the student model. Self-distillation (Gou et al., 2021) does not need an extra model as the teacher. Instead, intermediate representations in the model’s higher layers are taken as the teacher of the lower layers. This requires the higher layers’ representations to be better than the lower layers’. KD has not been fully explored in the OCL setting and SPOT (Kakogeorgiou et al., 2024) is the only one we know for now. But SPOT’s “self-distillation“ is actually offline distillation, which consumes at least two times training cost. Our DIAS realizes exact self-distillation by utilizing the definitively refined attention maps of the aggregation iterations, saving much cost.
@@ 3.1. Preliminary
where $i_{\mathrm{a}}$ is the total number of iterations for aggregation, and $i_{\mathrm{a}}=3$ for unconditional queries (Locatello et al., 2020) while $i_{\mathrm{a}}=1$ for conditional queries (Kipf et al., 2022). $\bm{A}^{i}_{\mathrm{a}}\in\mathbb{R}^{s\times h\times w}$ is aggregation attention map and can be converted to object segmentation with $\mathrm{argmax(\cdot)}$ along dimension $s$. $\bm{\phi}_{\mathrm{a}}$ can be parameterized into different Slot Attention variants like in Biza et al. (2023); Jia et al. (2023). We call $\bm{S}^{i}$ as “query slots” and $\bm{S}$ as “aggregated slots”. Note that $\bm{\phi}_{\mathrm{a}}$ enables all query vectors in $\bm{S}^{i}$ to compete with one another for some super-pixels in $\bm{Z}$.
@@
where $\bm{A}_{\mathrm{d}}\in\mathbb{R}^{s\times h^{\prime}\times w^{\prime}}$ is decoding attention map and can be converted to object segmentation masks by binarization; $\bm{Y}^{\prime}$ is the reconstructed $\bm{Y}$; and $\bm{Y}$ can be $\bm{X}$ (Locatello et al., 2020), $\bm{Z}$ (Seitzer et al., 2023), or $\bm{Z}$’s quantization (Zhao et al., 2025b), etc. $\bm{C}$ is the destructed $\bm{Y}$, which varies across methods.
@@ 3.2. Re-Initialized Aggregation
When there are fewer objects (plus the background) than the number of slots, redundant slots should be removed, or they will compete for object super-pixels and cause objects to be segmented into parts. SOLV (Aydemir et al., 2023) and AdaSlot (Fan et al., 2024) are the only two studies so far that explore this idea. They both conduct the redundancy reduction $\bm{\phi}_{\mathrm{r}}$ after the aggregation:
@@
Here Equation 8 is the initial aggregation, Equation 9 is the redundancy reduction, Equation 10 is the extra aggregation, and Equation 11 is the masked decoding. For the initial aggregation, we use $i_{\mathrm{a}}-1$ iterations, and $i_{\mathrm{a}}\geq 2$. For the redundancy reduction, we adopt SOLV’s solution, agglomerative clustering (Learn, 2025b) by cosine distance given some threshold. For the extra aggregation, we apply the redundancy mask $\bm{M}_{\mathrm{r}}$ to the query-key attention logits so as to discard the effect of those zeroed-out slots (Fan et al., 2024). For the masked decoding, we follow the solutions from SOLV and AdaSlot.
@@
69.4 ±8.7 89.9 ±1.3 59.4 ±2.9 57.8 ±2.4 22.1 ±0.8 42.7 ±0.2 32.8 ±0.1 31.3 ±0.4 27.3 ±4.6 37.7 ±2.5 46.1 ±2.1 45.0 ±1.7
Table 1. Object discovery on images. Using DINO2 ViT s/14 for OCL encoding; input resolution is 256$\times$256 (224$\times$224).

### v3

@@ Abstract.
Unlike popular solutions based on dense feature maps, Object-Centric Learning (OCL) represents visual scenes as sub-symbolic object-level feature vectors, termed slots, which are highly versatile for tasks involving visual modalities. OCL typically aggregates object superpixels into slots by iteratively applying competitive cross attention, known as Slot Attention, with the slots as the query. However, once initialized, these slots are reused naively, causing redundant slots to compete with informative ones for representing objects. This often results in objects being erroneously segmented into parts. Additionally, mainstream methods derive supervision signals solely from decoding slots into the input’s reconstruction, overlooking potential supervision based on internal information. To address these issues, we propose Slot Attention with re-Initialization and self-Distillation (DIAS): i) We reduce redundancy in the aggregated slots and re-initialize extra aggregation to update the remaining slots; ii) We drive the bad attention map at the first aggregation iteration to approximate the good at the last iteration to enable self-distillation. Experiments demonstrate that DIAS achieves state-of-the-art on OCL tasks like object discovery and recognition, while also improving advanced visual prediction and reasoning. Our source code and model checkpoints are available on https://github.com/Genera1Z/DIAS.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755339isbn: 979-8-4007-2035-2/2025/10submissionid: mfp3301copyright: noneccs: Computing methodologies Computer vision representations
@@
Figure 1. Our method is inspired by the following two key observations. (o1) We re-initialize an extra aggregation to update the remaining slots after slots redundancy reduction, instead of decoding the remaining slots directly. (o2) We self-distill the attention map at the first iteration to approximate the almost-always-good attention at the last aggregation iteration, rather than from aggregation attention to decoding attention.
@@ 1. Introduction
Humans’ vision system typically decomposes scenes into objects and allocates attention at the object level (chen2012objectbasedattention; kahneman1992objectfiles). This efficiently captures scene information, models object attributes and their interrelationships, and ultimately supports complex functionalities of prediction, reasoning, planning and decision-making (bar2004visual). Similarly, in Artificial Intelligence (AI), Object-Centric Learning (OCL) seeks to aggregate dense (super-)pixels of images or videos into sparse feature vectors, termed slots, representing corresponding objects in a sub-symbolic manner (locatello2020slotattent). In multimodal AI systems, object-centric representations can encode visual information more effectively and integrate with other modalities more flexibly than popular approaches based on dense feature maps (zhang2024omgllava; wang2024omnidrive).
@@
Existing OCL typically follows the encoding-aggregation-decoding paradigm (zhao2025vvo). Firstly, input image or video pixels are encoded into feature map; Secondly, object super-pixels in the feature map are iteratively aggregated into corresponding feature vectors, i.e., slots, using competitive cross attention, i.e., Slot Attention (locatello2020slotattent) or its variants (biza2023isa; jia2023boqsa); Lastly, these slots are decoded to reconstruct the input in some format, providing the self-supervision signal.
@@
However, (i1) once initialized, these slots are iterated naively in the aggregation, and when there are fewer objects than the predefined number of slots, redundant slots would compete with informative ones for object representation (aydemir2023solv; fan2024adaslot), causing objects to be segmented into parts. Additionally, (i2) mainstream methods (singh2021slate; seitzer2023dinosaur; wu2023slotdiffuz) derive supervision signals solely from reconstructing the input using the aggregated slots, leaving internal information unutilized, e.g., the aggregation attention maps at latter iterations have better object segmentation, which can be utilized to provide extra supervision signal.
@@
Issues (i1) and (i2) remain unresolved despite just a few ingenious attempts. (r1) SOLV (aydemir2023solv) and AdaSlot (fan2024adaslot) remove possible redundancy in the aggregated slots, but without further updating the remaining. As shown in Figure 1 o1, the remaining slots, despite being informative, are already distracted by redundant slots and are still of low quality in object representation. (r2) SPOT (kakogeorgiou2024spot) drives the aggregation attention map to approximate the whole decoding attention map, but without discarding the latter’s bad compositions. As shown in Figure 1 o2, the decoding attention map, despite being generally better, sometimes has lower accuracy in object segmentation than the aggregation attention map. This is why SPOT has to pretrain an extra copy of the model as the teacher and has to run both the teacher and student for their offline distillation, at least doubling the training cost.
@@
Our solution is simple yet effective. (s1) We use the remaining slots after slots redundancy reduction to re-initialize an extra aggregation. This improves slots’ object representation quality, instead of directly decoding the remaining slots. (s2) We use the attention map at the last aggregation iteration to guide that at the first iteration, where the guidance is mostly always better. This enables self-distillation, requiring no extra model copy, no pretraining, and minimal additional cost. Besides, (s3) we implement a generalized auto-regressive (AR) decoder for OCL. Since existing AR decoders’ fix-order flattening disrupts spatial correlations among superpixels (singh2021slate; singh2022steve; kakogeorgiou2024spot), we use random-order flattening to enforce the decoder to model spatial correlations.
@@
Figure 2. Our DIAS introduces three novel designs: (i) Reinitialized aggregation, which improves slots’ object representation quality by re-initializing an extra aggregation to update the remaining slots after slots redundancy reduction; (ii) Self-distilled aggregation, which obtains better internal supervision by approximating the almost-always-better attention map at the last aggregation iteration from that at the first aggregation iteration; and (iii) Random auto-regressive decoding, which enforces the decoder’s modeling of spatial correlations by randomly flattening a 2-dimensional feature map into a 1-dimensional sequence. On the top left is the common OCL architecture, which is adapted from (zhao2025vvo). $\bm{\phi}_{\mathrm{a}}$ is OCL aggregator, $\bm{\phi}_{\mathrm{d}}$ is OCL decoder, and $\bm{\phi}_{\mathrm{r}}$ is slots redundancy reduction. m is the mask token, bos is the Begin-of-Sentence token, and pos emb stands for the position embedding tensor. All notations are defined in the marked sections.
@@ 2. Relate Work
OCL aggregation. The core module of OCL is Slot Attention (locatello2020slotattent), an iterative cross attention that takes some feature vectors as the competitive query, with the input’s feature map as the key and value. BO-QSA (jia2023boqsa) improves the gradient flow on the query vectors for better optimization. ISA (biza2023isa) realizes scaling, rotation and translation invariance in slots with inductive biases. While others learn Gaussian distributions to sample the query vectors, SAVi (kipf2021savi) and SAVi++ (elsayed2022savipp) project conditions like object bounding boxes into query vectors. These aggregators are all faced with the issue that redundant slots compete with informative slots to represent single objects as parts. SOLV (aydemir2023solv) merges similar slots after aggregation using agglomerative clustering by some threshold. AdaSlot (fan2024adaslot) removes redundant slots after aggregation using a learned policy with some regularization. These two recent methods only remove or merge slots but do not update the remaining ones. Our DIAS improves the aggregation by both reducing redundant slots and updating the remaining slots with extra re-initialized aggregation.
@@
OCL decoding. There are mainly three types of decoding for OCL. Mixture-based decoders include CNN (kipf2021savi; elsayed2022savipp), MLP (seitzer2023dinosaur) and SlotMixer (zadaianchuk2024videosaur). The former two broadcast each of the aggregated slots into some spatial dimension, decode them respectively and mix the final components as the reconstruction. This means that the computation is proportional to the number of slots (seitzer2023dinosaur), which is very expensive. The latter one mixes the slots with a Transformer decoder without self-attention on the query. Such a decoder has to work in lower slot channel dimensions (zadaianchuk2024videosaur), which limits object representation quality. Auto-regression-based decoders include causal Transformer decoder (singh2021slate; singh2022steve) and causal Transformer decoder with nine permutations (kakogeorgiou2024spot). They basically drive the slots to reconstruct the input’s feature map with the causally masked input feature as the query, in a next-token prediction manner. However, flattening a 2-dimensional feature map into a 1-dimensional sequence in a fixed order loses the spatial relationship (kakogeorgiou2024spot). Diffusion-based decoders include different conditional Diffusion models (wu2023slotdiffuz; jiang2023lsd). They basically recover the noise from the noise diffused input with slots as conditions, to drive slots to be informative. But, Diffusion models are very large and sensitive to train (rombach2022latentdiffusion). Our DIAS realizes general auto-regression via arbitrary ordering. In terms of the reconstruction target, there are also works that introduce quantization (singh2021slate; singh2022steve; zhao2024gdr; zhao2024msf; zhao2025vvo), but we focus on the non-quantized case.
@@
Knowledge distillation. The classical offline knowledge distillation (KD) (gou2021kd) requires a pretrained model as the teacher and a to-be-trained model as the student. The teacher is frozen and runs at the same time as the student to provide better representations as targets for the corresponding student representations. There is also online distillation (gou2021kd) where no pretraining of a teacher model is needed. A teacher model typically has the same architecture as the student and its weights are usually updated by exponentially moving averaging (wiki2025ema) the student model. Self-distillation (gou2021kd) does not need an extra model as the teacher. Instead, intermediate representations in the model’s higher layers are taken as the teacher of the lower layers. This requires the higher layers’ representations to be better than the lower layers’. KD has not been fully explored in the OCL setting and SPOT (kakogeorgiou2024spot) is the only one we know for now. But SPOT’s “self-distillation“ is actually offline distillation, which consumes at least two times training cost. Our DIAS realizes exact self-distillation by utilizing the definitively refined attention maps of the aggregation iterations, saving much cost.
@@ 3.1. Preliminary
where $i_{\mathrm{a}}$ is the total number of iterations for aggregation, and $i_{\mathrm{a}}=3$ for unconditional queries (locatello2020slotattent) while $i_{\mathrm{a}}=1$ for conditional queries (kipf2021savi). $\bm{A}^{i}_{\mathrm{a}}\in\mathbb{R}^{s\times h\times w}$ is aggregation attention map and can be converted to object segmentation with $\mathrm{argmax(\cdot)}$ along dimension $s$. $\bm{\phi}_{\mathrm{a}}$ can be parameterized into different Slot Attention variants like in biza2023isa; jia2023boqsa. We call $\bm{S}^{i}$ as “query slots” and $\bm{S}$ as “aggregated slots”. Note that $\bm{\phi}_{\mathrm{a}}$ enables all query vectors in $\bm{S}^{i}$ to compete with one another for some super-pixels in $\bm{Z}$.
@@
where $\bm{A}_{\mathrm{d}}\in\mathbb{R}^{s\times h^{\prime}\times w^{\prime}}$ is decoding attention map and can be converted to object segmentation masks by binarization; $\bm{Y}^{\prime}$ is the reconstructed $\bm{Y}$; and $\bm{Y}$ can be $\bm{X}$ (locatello2020slotattent), $\bm{Z}$ (seitzer2023dinosaur), or $\bm{Z}$’s quantization (zhao2025vvo), etc. $\bm{C}$ is the destructed $\bm{Y}$, which varies across methods.
@@ 3.2. Re-Initialized Aggregation
When there are fewer objects (plus the background) than the number of slots, redundant slots should be removed, or they will compete for object super-pixels and cause objects to be segmented into parts. SOLV (aydemir2023solv) and AdaSlot (fan2024adaslot) are the only two studies so far that explore this idea. They both conduct the redundancy reduction $\bm{\phi}_{\mathrm{r}}$ after the aggregation:
@@
Here Equation 8 is the initial aggregation, Equation 9 is the redundancy reduction, Equation 10 is the extra aggregation, and Equation 11 is the masked decoding. For the initial aggregation, we use $i_{\mathrm{a}}-1$ iterations, and $i_{\mathrm{a}}\geq 2$. For the redundancy reduction, we adopt SOLV’s solution, agglomerative clustering (sklearn2025agglomerat) by cosine distance given some threshold. For the extra aggregation, we apply the redundancy mask $\bm{M}_{\mathrm{r}}$ to the query-key attention logits so as to discard the effect of those zeroed-out slots (fan2024adaslot). For the masked decoding, we follow the solutions from SOLV and AdaSlot.
@@
81.6 ±1.2 80.5 ±0.7 64.2 ±0.1 62.4 ±0.2 22.1 ±0.8 42.7 ±0.2 32.8 ±0.1 30.1 ±0.3 27.3 ±4.6 34.6 ±2.5 44.8 ±9.8 42.8 ±1.7
Table 1. Object discovery on images. Using DINO2 ViT s/14 for OCL encoding; input resolution is $\times256$ ($\times224$).